I haven't exhaustively looked at all cases, but it appears that, at least for the "bad data" case 05:10-06:10 AM on 5 June at Platteville, dew formation caused the bad h2o and presumably co2 values. During this case, RH is at least 97% and w'h2o' fluxes during the preceeding hour are negative (water deposition to the surface).
I note that this occurred despite the Li7500 being mounted at 45 degrees upside down, but presumably dew on the mirror also would affect the signal.
(This analysis was assisted by David Reed.)
RH.6m at Marshall (TRH010) looks wrong from Jun 9 06:10 to 07:30 MDT.
All the mobile ISS data were copied with rsync from the site back to the FTP data directory as of 2009-06-08 04:00 UTC.
iss1:/> cd /iss/ds iss1:/iss/ds> rsync -av ./ [email protected]:/net/ftp/pub/archive/iss/asp/iss1
This includes all the GAUS data, raw SPC, raw NIDAS, and NIMA SPC and netcdf.
8June09 ~10:00MDT
Platteville turned off, station returned to Boulder.
Since Jun 7 19:05 13:05 MDT Marshall adam was only recording the 2 inputs on the viper serial ports, not the emerald serial card. This is the PC104 interrupt issue that I thought I had fixed....
As a result we only have data from the 2m sonic (and GPS) from that time until the system was remotely soft-rebooted on Jun 8 09:04 MDT. It stayed up for 10 minutes and then hung. It was hard-booted at 10:00.
7June09 ~16:15MDT
Swapped Pocketec.
The area had some rain this afternoon and the Echo is reading ~34 now. Everything else seemed reasonable.
Data was uploaded in Boulder via porter.....
porter:/h/eol/militzer-> copy_pv_usbdisk.sh
usbdisk disk not mounted. Trying to mount /media/usbdisk...
source directory=/media/usbdisk/projects/ASP09
destination directory=/scr/isfs/projects
Filesystem 1K-blocks Used Available Use% Mounted on
/dev/sdd1 57685532 407292 54347988 1% /media/usbdisk
Filesystem 1K-blocks Used Available Use% Mounted on
eve.eol.ucar.edu:/Volumes/ISFS
1953415136 1534104160 419310976 79% /scr/isfs
/h/eol/militzer
sending incremental file list
ASP09/
rsync: failed to set times on "/scr/isfs/projects/ASP09": Permission denied (13)
ASP09/raw_data/
rsync: failed to set times on "/scr/isfs/projects/ASP09/raw_data": Permission denied (13)
ASP09/raw_data/platteville_20090606_221745.dat
16141416 100% 11.74MB/s 0:00:01 (xfer#1, to-check=2/5)
ASP09/raw_data/platteville_20090607_000000.dat
113755580 100% 10.60MB/s 0:00:10 (xfer#2, to-check=1/5)
ASP09/raw_data/platteville_20090607_120000.dat
98214290 100% 10.23MB/s 0:00:09 (xfer#3, to-check=0/5)
sent 228139407 bytes received 77 bytes 10139532.62 bytes/sec
total size is 228111286 speedup is 1.00
rsync error: some files/attrs were not transferred (see previous errors) (code 23) at main.c(1040) [sender=3.0.5]
Daily Report for Sat 6 June 2009
MISS at Marshall for ASP mini-field project
Fine, mainly sunny with light winds. Around mid-day on some mid-level clouds and widelyscattered low level CU. Also some high level cirrus. Less haze than yesterday. Brief brisk westerly breeze around 2330Z.
Four soundings 15Z, 18Z, 21Z and 0Z :
15Z was relaunched as first sonde string separated from RS92 unit.
21Z 7C temperature bias (high) on pre-launch checks. String didn't fully unwind so winds may be a little noisy.
0Z Lost GPS at launch. It was tracked visually to about 350mb flying in an almost striaght line at approx 60 deg azi and 60 deg elevation.
18Z and 21Z showed some evidense of gravity wave breaking in winds at upper levels.
Wind profiler mainly only seeing boundary layer scatter. Ceilometer seeing clouds around 4 - 5 km most of afternoon.
For two periods of an hour or more today, Rpile.in.mfs went to zero.
After trying to reseat connectors/wires in the logger box, the fix was simply to cut and restrip the wires inside the logger for the 2L/2H inputs.
6June09 ~16:15MDT
Swapped Pocketecs. Gordon merged.
Echo probe was reading ~ -67, tried the other probe more recently installed in the ground by CG, and it also was reading about the same. Left the original one connected to logger. Probably will be small spike in data from short swap.
Other data appeared reasonable from quick look.
State variables
---------------
P: ok
T: ok
RH: ok
Radiation
---------
Rnet: ok
Rsw: ok
Rlw: ok, but spike in Rlw.in between 0930 and 1000 on June 6
Rsum: see Rlw
Soil
Tsoil: ok
Qsoil: ok, but spike at 1520, June 5; swapped sensors?
Gsoil: ok
Csoil: see Qsoil
Prop Vane (pv)
---------
Spd: ok
Dir: ok
Sonics
------
spd: ok
dir: ok
w: ok
u'u': ok
v'v': ok
w'w': ok
u*: ok
tc: ok
tc'tc': ok
w'tc': ok
Licors
------
h2o: negative spike at pv 0600, June 5
h2o'h2o': positive spike at pv 0600, June 5
w'h2o': ok
co2: negative spike at pv 0600, June 5
opposite?? excursions at 4m and 10m at mfs; ~1800-1900, June 5
co2'co2': ok
w'co2': ok
Daily Report from MISS at Marshall in support of ASP mini-project
Fri June 5, 2009
Wave clouds over Bouler this morning, with increasing mid-level clouds during the day. Light southerly and easterly winds. Quite hazy. Late afternoon a thunderstorm with lots of lightning and oaccaionally heavy rain passed over.
Four soundings were launched : 1500 1730, 2030, 2330 UTC. All soundings went well with ASP students in attendance. Boundary layer appeared to be low (400 - 600m) in all cases. 1730 and 2030 showed evidence of gravity wave breaking at upper levels (above 100 mb). 2330 sonde was launched as the thunder storm approched and appeared to travel along with the storm to the NE.
Wind profiler worked well although not seeing much above the low boundary layer. During the early part of the storm there were cases of unusally clear separation of precip and clear-air signals in the Doppler spectra.
Surface station showed westerlies last night, not seen in the profiler so must have been low level, may be drainage flow out of Eldorado canyon.
Bill Brown (NCAR/EOL)
[Gary created this log entry from Bill's comment added to the wiki front page]
now is 1731 MDT. We had a trace of sprinkles perhaps 1.5 hours ago as well. This is enough to hear on the roof.
Rsw.out was clearly dirty (and appears to have a few specks of something inside the dome as well). Rnet might have had a bit of dirt as well. Accessed the sensors by standing on the (only) folding chair from the trailer. This done 1538-1544 starting with Rlw.out and ending with Rnet.
We've seen some spikes in the Marshall CO2 data at 10m (also some at 4m). We've also observed a raptor using the 10m boom as a perch. (I tried to take a few photos.) I guess it it possible that the Licor is measuring raptor breath! I update the bird time when I download my photos.